Suppose ∠A and ∠B are complementary angles. If ∠A = 2x-6 and ∠B = 3x+1, what is x and the measure of each angle?
Answer by LinnW(1048)

angle A = 2x-6
angle B = 3x+1
angle A + angle B = 90 since the angles are complementary
(2x-6) + (3x+1) = 90
5x - 5 = 90
add 5 to each side
5x = 95
divide each side by 5
x = 19
angle A = 2x-6
2(19)-6
38 - 6
32
angle B = 90 - 32
58
So angle A is 32 and angle B is 58.